OpenStudy (anonymous):

Which of the following is the slope between the points (-1, -5) and (7, -3)?

jimthompson5910 (jim_thompson5910):

Similar Example Find the slope of the line through the points (-3, 8) and (2, 7) m = (y2-y1)/(x2-x1) m = (7-8)/(2-(-3)) m = (7-8)/(2+3) m = -1/5 So the slope of the line that passes through the two points (-3, 8) and (2, 7) is -1/5 Note: this is NOT the answer as it's just an example

OpenStudy (anonymous):

but when i did it i got -8/8 and thats not a choice. even after i simplified it

jimthompson5910 (jim_thompson5910):

(x1,y1) = (-1, -5) ---> x1 = -1, y1 = -5 (x2,y2) = (7, -3) ---> x2 = 7, y2 = -3

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (-3 - (-5))/(7 - (-1)) m = (-3 + 5)/(7 + 1) m = ??

OpenStudy (anonymous):

i have to be wrong. i got 2/8

jimthompson5910 (jim_thompson5910):

2/8 reduces to ???

OpenStudy (anonymous):

1/4?

jimthompson5910 (jim_thompson5910):

always make sure to reduce as much as possible

jimthompson5910 (jim_thompson5910):

yes good

OpenStudy (anonymous):

yay! thank you!!

jimthompson5910 (jim_thompson5910):

yw